# AgentsCoMerge: Large Language Model Empowered Collaborative Decision Making for Multi-Lane Merging

### v1

@@ 2.3 Multi-Agent Collaborative Decision-Making
In the domain of multi-agent collaboration and communication, there has been a notable interest in utilizing LLMs for enhancing cooperative efforts [25, 26, 27, 7, 28, 29, 30]. Zhang et al. [31] designed a system utilizing LLMs to enhance cooperative behavior among embodied agents, facilitating effective communication to achieve common objectives in diverse settings. Hong et al. [32] created MetaGPT to incorporate human involvement in LLM-guided collaborative activities, thus combining human intuition with the capabilities of artificial intelligence. Furthermore, Li et al. [33] introduced CAMEL, a system that establishes role-based agent interactions, fostering autonomous collaboration through dynamic role engagement. Specifically, in the realm of connected and autonomous vehicles (CAVs), the emphasis on collaborative decision-making utilizing LLMs has intensified in accident analysis, scene simulation, and intersection management. For instance, Wang et al. [34] developed AccidentGPT, an LLM-based framework integrating multi-sensor data to facilitate comprehensive accident analysis and prevention for traffic safety. Subsequently, Wei et al. [35] introduced ChatSim, an LLM-driven simulation tool for AD scenarios, enabling highly flexible scene editing through agent collaboration. Most recently, Hu et al. [7] developed a collaborative driving framework employing LLMs to accumulate knowledge and experience for lifelong learning, which demonstrated its superior communication and collaboration efficiency in highway and intersection scenarios. However, in our paper, we focus on harnessing LLMs and VLMs to facilitate comprehensive collaborative decision-making under on-ramp merging scenarios. By leveraging the advanced capabilities of foundation models, we aim to develop a system to coordinate and manage decision-making processes among multiple agents under multi-lane merging scenarios, thereby advancing the safety, efficiency, and reliability of connected and autonomous driving (CAD).
@@ 3.2 Vehicle Formulation
$\dot{x}$ $=u\cos\alpha,$ (1)
$\dot{y}$ $=u\sin\alpha.$ (2)
The non-holonomic constraint for the front wheel is
@@
Then, the kinematic constraint for the front wheel can be rewritten as
$\dot{x}_{f}\sin(\alpha+\beta)-\dot{y}_{f}\cos(\alpha+\beta)-\dot{\alpha}R\cos\beta=0.$ (6)
From Eq. (1), (2) and (6), we can derive
@@
$s=\varphi(y,\dot{y},\ldots,y^{(b)}),$
$c=\gamma(y,\dot{y},\ldots,y^{(b)}).$
@@
The $\dot{\alpha}$ can be obtained by differentiating Eq. (10),
$\dot{\alpha}=\frac{\dot{\delta_{1}}\ddot{\delta_{2}}-\ddot{\delta_{1}}\dot{\delta_{2}}}{\dot{\delta_{1}}^{2}+\dot{\delta_{2}}^{2}}.$ (11)
Plugging Eq. (11) into Eq. (7), we can obtain the steering angle $\beta$ as
$\beta=\arctan\frac{(\dot{\delta_{1}}\ddot{\delta_{2}}-\dot{\delta_{2}}\ddot{\delta_{1}})R}{\left({\dot{\delta_{1}}^{2}+\dot{\delta_{2}}^{2}}\right)^{\frac{3}{2}}}.$ (12)
@@
From the above derivations of the formulation, we can see that given the output $\delta=(x,y)$, which is three times differentiable, we can express other states of the vehicle $[\alpha,\beta]^{T}$ and the control inputs $[u,\omega]^{T}$ purely in terms of the $\delta$ and its derivatives.
@@ 3.3 Collaborative Decision-Making Formulation
Multi-vehicle collaborative decision-making can be modeled as a Decentralized Partially Observable Markov Decision Process (DPOMDP) with inter-agent communication (DPOMDP-Com).
A DPOMDP is a sequential decision-making problem involving multiple agents (e.g., CAVs) in an environment where information is incomplete and uncertain. Each agent independently takes actions based on its own observations. These actions are taken simultaneously and rely solely on the agent’s local observation. The outcomes of these actions provide immediate rewards and generate new observations for each agent, guided by probabilistic models that follow Markov properties and span continuous states, actions, and observation spaces. Mathematically, the DPOMDP model can be described as a tuple:
@@
Furthermore, the communication among agents can be incorporated into the observations and actions of the agents. If the communication is free, instantaneous and lossless, the DPOMDP model can be simplified to POMDP, which means all agents can get all observations with global information. If the communication has a delay or cost, the agents should decide when to communicate and what to communicate. In this case, we can reformulate the DPOMDP model as a DPOMDP with inter-agent communication (DPOMDP-Com), which contains a set of messages $\mathbf{M}$. The complexity of DPOMDP-Com is the same as DPOMDP, but it will be beneficial for decision-making to consider inter-agent communication.
@@ 4 Method
Figure 3: Overall Architecture of AgentsCoMerge. The overall architecture of the proposed collaborative driving framework under multi-lane merging scenarios.
@@ 4.1 Overview
Fig. 3 shows our multi-vehicle, closed-loop, and lifelong learning collaborative decision-making framework under multi-lane merging scenarios. It contains four main components: 1) scene observation and understanding (Section 4.2), 2) hierarchical planning (Section 4.3), and 3) inter-agent communication (Section 4.4). The scene observation and understanding module is responsible for perceiving the surrounding environments and extracting necessary information. The hierarchical planning module is designed to perform multi-round reasoning based on the inputs and generate the final decision. The inter-agent communication module is responsible for enabling the agents to communicate with each other. In addition, we propose a reinforcement reflection guided training paradigm (Section 4.5) to enhance the capabilities of reasoning and decision-making. which consists of two steps: 1) planning fine-tuning and 2) reinforcement reflection fine-tuning. The details about each component are discussed in the following sections.
@@ 4.2.1 Vision-based Scene Understanding
After obtaining the visual features $\mathbf{F}_{\text{vis}}$, we then feed them into the token alignment module $f_{\text{align}}(\cdot)$ to align the visual tokens to the space of text tokens. Specifically, we design a single-layer cross-attention module that is initialized with random values. It employs a set of trainable vectors (Embeddings) $\mathbf{Q_{\text{vis}}}$ as query vectors, while using the visual features $\mathbf{F}_{\text{vis}}$ from the visual encoder as keys and the values in the cross-attention operations. Then we can obtain the cross attention matrix $\mathbf{\Phi}_{\text{vis}}$:
@@ 4.3 Hierarchical Planning
Reasoning Process. In order to obtain decisions, an agent needs to reason over all the information. Given the complexity and diversity of the information, a robust reasoning engine is crucial for enabling the agent to make effective and accurate decisions. To enhance reasoning capabilities, we directly employ LLM as the reasoning engine with carefully designed prompts to process all the information and generate the decision.
@@
Trajectory Planning. After obtaining the meta-action decision, the agent needs to refine the decision into a specific trajectory $\mathcal{T}=\{(x_{1},y_{1}),(x_{2},y_{2}),\ldots,(x_{t},y_{t})\}$. Note that we can directly represent the trajectory as a sequence of words:
@@
where $\text{Tokenizer}(\cdot)$ is a function that converts the trajectory into a sequence of words and $w_{i}$ is the $i$-th word in the sequence. With this language representation, we can reformulate the trajectory planning problem as a language generation problem:
@@
where $w$ and $\hat{w}$ represent the words from the planned trajectory $\mathcal{T}$ and the human driving trajectory $\hat{\mathcal{T}}$, respectively. By focusing on increasing the probability $P$ of the words $\hat{w}$ derived from the human driving trajectory $\hat{\mathcal{T}}$, motion planners can create driving trajectories that more closely mimic human driving behavior. In addition, we can reformulate the language generation problem as an optimization problem:
@@ 4.5 Reinforcement Reflection Guided Training
In the second step, we utilize detailed textual reflection about errors in meta-actions and waypoint prediction to enhance the learning process of the agent. This approach allows the agent to gain an impression from successes and failures, which can offer a particularly instructive form of supervision. Specifically, we first define the failure cases including collisions, deviation of route, low efficiency, and low comfort, if these failure cases occur, we will prompt the agent to reflect on its decisions and predictions. The example of this process is shown in Fig. 6. The agent is then trained to minimize the loss function, which is defined as the weighted sum over the language cross-entropy loss and the waypoint prediction mean square error (MSE) loss:
@@ 5.1 Benchmarks and Metrics
Closed-loop Benchmarks. In the closed-loop evaluation, we leverage the LimSim++ simulator [57]. LimSim++ is a platform that can simulate the fine-grained dynamic environment and can provide a continuous simulation under different traffic scenarios. In addition, it can allow agents to interact with different vehicles. In order to evaluate the performance of our method in this simulator, we leverage the following metrics: Comfort Score (CS), Efficiency Score (ES), Safety Score (SS), and Driving Score (DS). CS measures the comfort of the driving behavior and can be formulated as
$\text{CS}=(s_{x_{a}}+s_{y_{a}}+s_{x_{j}}+s_{y_{j}})/4,$ (27)
where $s_{x_{a}}$ and $s_{y_{a}}$ are the reference values of lateral and longitudinal acceleration of an agent, and $s_{x_{j}}$ and $s_{y_{j}}$ are the reference values of lateral and longitudinal jerk of an agent. ES measures the efficiency of the driving behavior and can be formulated as
$\text{ES}=\left\{\begin{aligned} &1.0,&\text{if}\ v_{\text{ego}}\geq v_{0}\\ &v_{\text{ego}}/v_{0},&\text{else}\end{aligned}\right.,$ (28)
where $v_{\text{ego}}$ is the velocity of an agent and $v_{0}\in\{v_{\text{avg}},v_{\text{limt}}\}$. $v_{\text{avg}}$ is the average velocity of the surrounding vehicles and $v_{\text{limt}}$ is the speed limit. SS measures the safety of the driving behavior assessed by time to conflict (TTC), because if TTC is less than a specific threshold, it means that the agent is in a dangerous situation. SS can be formulated as
@@
where $t_{\text{ego}}$ is the TTC of an agent and $t_{t}$ is the threshold. Finally, DS measures the overall driving behavior and can be formulated as
@@
where $\alpha$ and $\beta$ are the penalty factors for collisions and speed violations, respectively. $\lambda_{1}$ and $\lambda_{2}$ represent the times at which three different types of punishable behaviors occur throughout the trajectory. $k_{1}$, $k_{2}$, and $k_{3}$ are the weights of CS, ES, and SS, respectively.
@@
Open-loop Benchmarks. In the open-loop evaluation, we leverage the nuScenes [58] dataset. The nuScenes dataset is a large-scale dataset for autonomous driving, which contains 1000 scenes, each of which contains 20 seconds of data. In order to evaluate our method in this dataset, we leverage L2 error and the collision rate. The L2 error measures the distance between the predicted trajectory and the ground truth trajectory, which can be formulated as
@@
The collision rate can be expressed as the ratio of the number of times a planned trajectory collides with other objects in a scene to the duration of the planned trajectory. The collision rate can be formulated as

### v2

# AgentsCoMerge: Large Language Model Empowered Collaborative Decision Making for Ramp Merging

@@ 2.3 Multi-Agent Collaborative Decision-Making
In the domain of multi-agent collaboration and communication, there has been a notable interest in utilizing LLMs for enhancing cooperative efforts [25, 26, 27, 7, 28]. Zhang et al. [29] designed a system utilizing LLMs to enhance cooperative behavior among embodied agents, facilitating effective communication to achieve common objectives in diverse settings. Hong et al. [30] created MetaGPT to incorporate human involvement in LLM-guided collaborative activities, thus combining human intuition with the capabilities of artificial intelligence. Furthermore, Li et al. [31] introduced CAMEL, a system that establishes role-based agent interactions, fostering autonomous collaboration through dynamic role engagement. Specifically, in the realm of connected and autonomous vehicles (CAVs), the emphasis on collaborative decision-making utilizing LLMs has intensified in accident analysis, scene simulation, and intersection management. For instance, Wang et al. [32] developed AccidentGPT, an LLM-based framework integrating multi-sensor data to facilitate comprehensive accident analysis and prevention for traffic safety. Most recently, Hu et al. [7] developed a collaborative driving framework employing LLMs to accumulate knowledge and experience for lifelong learning, which demonstrated its superior communication and collaboration efficiency in highway and intersection scenarios. However, in our paper, we focus on harnessing LLMs and VLMs to facilitate comprehensive collaborative decision-making under on-ramp merging scenarios. By leveraging the advanced capabilities of foundation models, we aim to develop a system to coordinate and manage decision-making processes among multiple agents under multi-lane merging scenarios, thereby advancing the safety, efficiency, and reliability of connected and autonomous driving (CAD).
@@ 3.2 Vehicle Formulation
$\dot{x}_{f}$ $=u\cos\alpha,$ (1)
$\dot{y}_{f}$ $=u\sin\alpha.$ (2)
where $\dot{x}_{f}$ and $\dot{y}_{f}$ are the velocities of the front wheel. The overdot means the derivative of the variable with respect to time. In addition, the front wheel’s velocity must align with its steering direction $\alpha+\beta$, the non-holonomic constraint for the front wheel is
@@
Differentiating $x_{f}$ and $y_{f}$ with respect to time, we obtain
$\dot{x}_{f}$ $=\dot{x}-R\sin\alpha\cdot\dot{\alpha},$ (6)
$\dot{y}_{f}$ $=\dot{y}+R\cos\alpha\cdot\dot{\alpha}.$ (7)
Substituting $\dot{x}=u\cos\alpha$ and $\dot{y}=u\sin\alpha$, we obtain
$\dot{x}_{f}$ $=u\cos\alpha-R\sin\alpha\cdot\dot{\alpha},$ (8)
$\dot{y}_{f}$ $=u\sin\alpha+R\cos\alpha\cdot\dot{\alpha}.$ (9)
Then, we can substitute $\dot{x}_{f}$ and $\dot{y}_{f}$ into Eq. (3), we have
$\left(u\cos\alpha-R\sin\alpha\cdot\dot{\alpha}\right)\sin(\alpha+\beta)=$
$\left(u\sin\alpha+R\cos\alpha\cdot\dot{\alpha}\right)\cos(\alpha+\beta).$
Thus, we can expand and group terms involving $u$ and $\dot{\alpha}$ to arrive at
$u\Big{[}\cos\alpha\sin(\alpha+\beta)-\sin\alpha\cos(\alpha+\beta)\Big{]}$displaystyle--
$R\dot{\alpha}\Big{[}\sin\alpha\sin(\alpha+\beta)+\cos\alpha\cos(\alpha+\beta)\Big{]}$ $=0,$
which can be simplified to
$u\sin\beta-R\dot{\alpha}\cos\beta=0.$ (10)
Therefore, we finally have
@@
$s$ $=\varphi(y,\dot{y},\ldots,y^{(b)}),$ (13)
$c$ $=\gamma(y,\dot{y},\ldots,y^{(b)}).$ (14)
@@
$\dot{\alpha}$ can be obtained by differentiating Eq. (16),
$\dot{\alpha}=\frac{d}{dt}\arctan\left(\frac{\dot{\delta}_{2}}{\dot{\delta}_{1}}\right)=\frac{\dot{\delta}_{1}\ddot{\delta}_{2}-\ddot{\delta}_{1}\dot{\delta}_{2}}{\dot{\delta}_{1}^{2}+\dot{\delta}_{2}^{2}}.$ (17)
From the kinematic model, we have $\dot{\alpha}=\frac{u}{R}\tan\beta$. With $\dot{\alpha}$ and $u=\sqrt{\dot{\delta}_{1}^{2}+\dot{\delta}_{2}^{2}}$, we obtain
$\beta=\arctan\left(\frac{\dot{\alpha}R}{u}\right)=\arctan\left(\frac{(\dot{\delta}_{1}\ddot{\delta}_{2}-\dot{\delta}_{2}\ddot{\delta}_{1})R}{(\dot{\delta}_{1}^{2}+\dot{\delta}_{2}^{2})^{3/2}}\right).$ (18)
@@
From the above derivations, we observe that given the output $\delta=(x,y)$, which is three times differentiable, we can express other states of the vehicle $[\alpha,\beta]^{T}$ and the control inputs $[u,\omega]^{T}$ purely in terms of the $\delta$ and its derivatives. This feature avoids solving differential equations during planning, enabling real-time applications in CAVs [43].
@@ 3.3 Collaborative Decision-Making Formulation
Multi-vehicle collaborative decision-making can be modeled as a Decentralized Partially Observable Markov Decision Process (DPOMDP) [44] with inter-agent communication (DPOMDP-Com) [45].
A DPOMDP is a sequential decision-making problem involving multiple agents (e.g., CAVs) in an environment where information is incomplete and uncertain. Each agent independently takes actions based on its own observations. These actions are taken simultaneously and rely solely on the agent’s local observation. The outcomes of these actions provide immediate rewards and generate new observations for each agent, guided by probabilistic models that follow Markov properties and span continuous states, actions, and observation spaces. Mathematically, the DPOMDP model can be described as a tuple [45]:
@@
Furthermore, the communication among agents can be incorporated into the observations and actions of the agents. If the communication is free, instantaneous and lossless, the DPOMDP model can be simplified to POMDP, which means all agents can get all observations with global information. If the communication has a delay or cost, the agents should decide when to communicate and what to communicate. In this case, we can reformulate the DPOMDP model as a DPOMDP with inter-agent communication (DPOMDP-Com), which contains a set of messages $\mathbf{M}$ [45]. The complexity of DPOMDP-Com is the same as DPOMDP, but it will be beneficial for decision-making to consider inter-agent communication.
@@ 4 Method
Figure 3: Overall Architecture of AgentsCoMerge. The overall architecture of the proposed collaborative driving framework under multi-lane merging scenarios. It includes four parts: 1) scene understanding, 2) planning, 3) inter-agent communication, and 4) reinforcement reflection. First, the agent understands the scene and environments by scene observation and understanding module. Then, the agent plans and makes decisions based on the environments and messages from other agents. Finally, the agent reflects on the decisions and improves its decision-making capability with reinforcement reflection guided training.
@@ 4.1 Overview
Fig. 3 shows our multi-vehicle, closed-loop, and lifelong learning collaborative decision-making framework under multi-lane merging scenarios. It contains four main components: 1) scene observation and understanding (Section 4.2), 2) hierarchical planning (Section 4.3), and 3) inter-agent communication (Section 4.4). The scene observation and understanding module is responsible for perceiving the surrounding environments and extracting necessary information. The hierarchical planning module is designed to perform multi-round reasoning based on the inputs and generate the final decision. The inter-agent communication module is responsible for enabling the agents to communicate with each other. In addition, we propose a reinforcement reflection guided training paradigm (Section 4.5) to enhance the capabilities of reasoning and decision-making, which consists of two steps: 1) planning fine-tuning and 2) reinforcement reflection fine-tuning. The details about each component are discussed in the following sections.
Algorithm 1 Pseudo-code of AgentsCoMerge Framework
1:Initialize:
2:$\mathbf{A}=\{a_{i}\}_{i=1}^{N}$triangleright CAV agent set
[…]
4:for $e=1\ \text{to}\ E_{\text{max}}$ do
5:     $\mathbf{o}_{0}\leftarrow\mathbf{S}.\text{reset}()$triangleright Initial observation
6:     for $t=1\ \text{to}\ T$ do
7:         Scene Observation & Understanding:
8:         $\mathbf{s}_{t}\leftarrow\mathbf{S}.\text{get\_sensor\_data}(a_{i})$
9:         $\mathbf{\Phi}_{\text{vis}}\leftarrow f_{\text{align}}(f_{\text{encode}}(\mathbf{s}_{t}))$triangleright Visual tokens
10:         $\boldsymbol{\tau}_{t}\leftarrow g_{\text{text}}(\mathbf{s}_{t})$triangleright Text tokens
11:         $\mathbf{r}_{i,t}\leftarrow h_{\text{scene}}(\boldsymbol{\tau}_{t},\mathbf{\Phi}_{\text{vis}})$triangleright Scene representation
12:         Hierarchical Planning:
13:         $\mathcal{M}_{i,t}\leftarrow\{\mathbf{m}_{j\rightarrow i}\}_{j\in\mathcal{N}_{i}}$triangleright Received messages
14:         $\boldsymbol{\phi}_{i,t}\leftarrow h_{\text{LLM}}(\{\mathbf{r}_{i,t}\}\cup\mathcal{M}_{i,t})$triangleright Reasoning
15:         $(\alpha_{i,t},\mathcal{T}_{i,t})\leftarrow\psi(\boldsymbol{\phi}_{i,t})$triangleright Meta-action & Traj.
16:         $\mathbf{x}_{i,t+1}\leftarrow f_{\text{update}}(\alpha_{i,t},\mathcal{T}_{i,t})$
17:         Inter-Agent Communication:
18:         if $a_{i}\in\mathcal{C}$ then $\triangleright$ $\mathcal{C}$: collaborative area
19:              $\mathbf{m}_{i,t}\leftarrow h_{\text{LLM}}(\alpha_{i,t})$triangleright Generate message
20:              $\text{Broadcast}(a_{i},\mathbf{m}_{i,t})$
21:         end if
22:         Execute & Observe:
23:         $\mathbf{S}.\text{execute}(\alpha_{i,t},\{\mathcal{T}_{i,t}\}_{i=1}^{N})$
24:         $[\text{failures}]\leftarrow\mathbf{S}.\text{step}()$triangleright Failure cases
25:         Reinforcement Reflection:
[…]
27:              $\boldsymbol{\rho}_{i,t}\leftarrow h_{\text{LLM}}(\mathcal{T}_{i,t},c_{t})$triangleright Reflection
28:              $\mathcal{L}_{\text{reflection}}=\mathcal{L}_{\text{LM}}(\boldsymbol{\rho}_{i,t},\mathbf{y}_{t})+\alpha\|\mathcal{T}_{i,t}-\hat{\mathcal{T}}_{i,t}\|_{2}^{2}$
29:              $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}_{\text{reflection}}$triangleright LoRA update
30:         end if
31:     end for
32:end for
@@ 4.2.1 Vision-based Scene Understanding
After obtaining the visual features $\mathbf{F}_{\text{vis}}$, we then feed them into the token alignment module $f_{\text{align}}(\cdot)$ to align the visual tokens to the space of text tokens. Specifically, we design a single-layer cross-attention module that is initialized with random values. It employs a set of trainable vectors (Embeddings) $\mathbf{Q_{\text{vis}}}$ as query vectors, while using the visual features $\mathbf{F}_{\text{vis}}$ from the visual encoder as keys and the values in the cross-attention operations. Then we can obtain the cross attention matrix $\mathbf{\Phi}_{\text{vis}}$ [10]:
@@ 4.3 Hierarchical Planning
Reasoning Process. In order to obtain decisions, an agent needs to reason over all the information. Given the complexity and diversity of the information, a robust reasoning engine is crucial for enabling the agent to make effective and accurate decisions. To enhance reasoning capabilities, we directly employ LLM as the reasoning engine with carefully designed prompts combined with the observation tokens (visual tokens and text tokens) to process all the information and generate the decision.
@@
Trajectory Planning. After obtaining the meta-action decision, the agent needs to refine the decision into a specific trajectory $\mathcal{T}=\{(x_{1},y_{1}),(x_{2},y_{2}),\ldots,(x_{t},y_{t})\}$. Note that we can directly represent the trajectory as a sequence of words [48]:
@@
where $\text{Tokenizer}(\cdot)$ is a function that converts the trajectory into a sequence of words and $w_{i}$ is the $i$-th word in the sequence. With this language representation, we can reformulate the trajectory planning problem as a language generation problem [48]:
@@
where $w$ and $\hat{w}$ represent the words from the planned trajectory $\mathcal{T}$ and the input tokens (history trajectory, scene understanding, etc.), respectively. By focusing on increasing the probability $P$ of the words $\hat{w}$ derivable from the input tokens, the LLM can create driving trajectories more precisely. In addition, we can reformulate the language generation problem as an optimization problem:
@@
Output of hierarchical planning. The output of the hierarchical planning is the meta-action decision and the trajectory, as well as the reasoning process.
@@ 4.5 Reinforcement Reflection Guided Training
In the second step, we utilize detailed textual reflection about errors in meta-actions and waypoint prediction to enhance the learning process of the agent. This approach allows the agent to gain an impression from successes and failures, which can offer a particularly instructive form of supervision. Specifically, we first define the failure cases including collisions, deviation of route, low efficiency, and low comfort, if these failure cases occur, we will prompt the agent to reflect on its decisions and predictions. The example of this process is shown in Fig. 6. The failure cases are described as follows:
  • Collision: To detect potential collisions, we use the kinematic bicycle model ${K}_{veh}$ to predict the future states $\mathbf{S}^{i}_{t+1}=(x_{t+1}^{i},y_{t+1}^{i},\alpha_{t+1}^{i},\beta_{t+1}^{i})$ of surrounding vehicles based on their current states $\mathbf{S}^{i}_{t}=(x_{t}^{i},y_{t}^{i},\alpha_{t}^{i},\beta_{t}^{i})$ and control inputs $\mathbf{C}_{t}^{i}=[u_{t}^{i},w_{t}^{i}]^{T}$. We then compute the intersections over union (IoU) between the ego vehicle’s bounding box $B_{t}^{s}$ and other vehicles’ bounding boxes $O_{t}^{i}$ at each predicted waypoint. A collision is detected if $\text{IoU}(B_{t}^{s},O_{t}^{i})>\varepsilon_{col}$, where $\varepsilon_{col}$ is a predefined collision threshold.
[…]
  • Low Efficiency: We evaluate the driving efficiency using a score Efficiency Score (ES) that compares the ego vehicle’s velocity $v_{ego}$ with a reference velocity $v_{0}$, which is either the average velocity of surrounding vehicles $v_{avg}$ or the speed limit $v_{lmt}$. The efficiency score is calculated as follows
    $\text{ES}=\begin{cases}1.0,&\text{if }v_{ego}\geq v_{0}\\ v_{ego}/v_{0},&\text{else}\end{cases}$ (30)
    where $v_{0}=\min\{v_{avg},v_{lmt}\}+\sigma$. $\sigma$ is the standard deviation of $v_{0}$, we use $\sigma=0$ in this study and can choose $2\sigma$ or $3\sigma$ according to the past observations. A low efficiency state is detected when $\text{ES}<\varepsilon_{E}$ for the threshold $\varepsilon_{E}$.
  • Low Comfort: The comfort of driving behavior is measured using a Comfort Score (CS) that considers both lateral and longitudinal aspects of motion:
    $\text{CS}=(s_{x_{a}}+s_{y_{a}}+s_{x_{j}}+s_{y_{j}})/4$ (31)
    where $s_{x_{a}}$ and $s_{y_{a}}$ are the reference values of lateral and longitudinal acceleration, respectively, and $s_{x_{j}}$ and $s_{y_{j}}$ are the reference values of lateral and longitudinal jerk. A low comfort state is triggered when $\text{CS}<\varepsilon_{C}$ for the threshold $\varepsilon_{C}$.
The agent is then trained to minimize the loss function, defined as the weighted sum over the language cross-entropy loss and the waypoint prediction mean square error (MSE) loss:
@@ 5.1 Benchmarks and Metrics
Closed-loop Benchmarks. In the closed-loop evaluation, we leverage the LimSim++ simulator [55]. LimSim++ is a platform that can simulate the fine-grained dynamic environment and can provide a continuous simulation under different traffic scenarios. In addition, it can allow agents to interact with different vehicles.
In order to evaluate the performance of our method in this simulator, we leverage the following metrics: Comfort Score (CS), Efficiency Score (ES), Safety Score (SS), and Driving Score (DS) [55]. CS measures the comfort of the driving behavior and can be formulated as Eq. 31. ES measures the efficiency of the driving behavior and can be formulated as Eq. 30. SS measures the safety of the driving behavior assessed by time to conflict (TTC), which estimates the time remaining before a potential collision occurs between two agents (e.g., vehicles, pedestrians) if both maintain their current velocity and trajectory. TTC is calculated by ${d}/{v_{\text{rel}}}$, where $d$ is the current distance between the ego vehicle and the object ahead, $v_{\text{rel}}=v_{\text{ego}}-v_{\text{lead}}$ is relative speed. If $v_{\text{rel}}\leq 0$, TTC is considered infinite (no collision under current conditions). Therefore, $\text{TTC}\in[0,\infty)$.
If TTC is less than a specific threshold, it means that the agent is in a dangerous situation. SS can be formulated as
@@
where $t_{\text{ego}}\in[0,\infty)$ is the TTC of an agent and $t_{t}$ is the threshold. Finally, DS measures the overall driving behavior and can be formulated as
@@
where $\alpha\in[0,1)$ and $\beta\in[0,1)$ are the penalty factors for collisions and speed violations (speed violation is triggered when the speed exceeds 11.11m/s), respectively. $\lambda_{1}$ and $\lambda_{2}$ represent the times at which two different types of punishable behaviors occur throughout the trajectory. $k_{1}$, $k_{2}$, and $k_{3}$ are the weights of CS, ES, and SS, respectively.
@@
Open-loop Benchmarks. In the open-loop evaluation, we leverage the nuScenes [56] dataset. The nuScenes dataset is a large-scale dataset for autonomous driving, which contains 1000 scenes, each of which contains 20 seconds of data. In order to evaluate our method in this dataset, we leverage L2 error and the collision rate. The L2 error measures the distance between the predicted trajectory and the ground truth trajectory, which is formulated as follows [57].
@@
The collision rate can be expressed as the ratio of the number of times a planned trajectory collides with other objects in a scene to the duration of the planned trajectory. The collision rate is formulated as follows [57].
@@
In addition, we train and evaluate the performance on HighD dataset [58]. The HighD dataset is a naturalistic vehicle trajectory dataset recorded from drone videos on German highways. It consists of 60 video recordings (instances), covering a cumulative driving duration of approximately 16.5 hours. The dataset involves over 110,500 vehicles, totaling around 45,000 km of driving data, and it captures approximately 5,600 lane-change maneuvers. The dataset was collected at six distinct highway sections (stretches), enabling evaluation across various road conditions. In order to learn the HighD dataset, we split the dataset into training set (the first 50 recordings) and testing set (the last 10 recordings). Then, we extract the trajectories into 10s and split them into 6s of past and 4s of the future trajectories pairs, which are used to train or evaluate the model. The 6s of past trajectories are feed into the model to predict the 4s of the future trajectories. In order to calculate the loss, we did not use the traditional RMSE loss function, because in our design, the trajectories are generated by the LLM, and we reformulate the trajectory planning problem as a language generation problem as stated in Section 4.3. The objective function is formulated as Eq. 29. In order to evaluate the performance on the HighD dataset, we use the root mean square error (RMSE) as evaluation metrics. The RMSE can measure the average deviation between the predicted and the actual values. It is calculated as:
$\text{RMSE}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}(y_{i}-\hat{y}_{i})^{2}}$ (39)
where $N$ is the number of samples, $y_{i}$ is the actual value, and $\hat{y}_{i}$ is the predicted value.
TABLE V: Performance Evaluation on HighD Dataset
Method RMSE in Prediction Horizons (s) Avg.
1s 2s 3s 4s
CS-LSTM [59] 0.61 1.27 2.09 3.10 1.77
Dual-Trans [60] 0.41 0.79 1.11 1.40 0.93
WSiP [61] 0.20 0.60 1.21 2.07 1.02
Social GAN [62] 0.30 0.78 1.46 2.34 1.22
Ours 0.25 0.53 0.91 1.22 0.73